OpenStudy (anonymous):

Find an equation for the line tangent to the graph of f(x)=sqrt(x)/5x+5 at the point (3,f(3))

OpenStudy (anonymous):

find \(f'(3)\) that is your slope then your point is \((3,f(3))\) point slope formula will give you the equation for the line

OpenStudy (anonymous):

it will be nice and ugly since \(f(3)=\frac{\sqrt{3}}{20}\) if i am reading your function correctly

OpenStudy (anonymous):

-1/80sqrt(3)?

OpenStudy (anonymous):

is it \(f(x)=\frac{\sqrt{x}}{5x+5}\) ?

OpenStudy (anonymous):

yes.. i got that ^ for f'(3)

OpenStudy (anonymous):

(1-3)/(10 sqrt(3) (1+3)^2)

OpenStudy (anonymous):

yeah i guess that is right http://www.wolframalpha.com/input/?i=+%281-3%29%2F%2810+sqrt%283%29+%283%2B1%29^2%29

OpenStudy (anonymous):

that's the slope?

OpenStudy (anonymous):

yup ugly huh?

OpenStudy (anonymous):

sqrt(3)/20=(-1)/(80sqrt(3))(3)+b?

OpenStudy (anonymous):

got it! thanks!

OpenStudy (anonymous):

you need an \(x\) and a \(y\) in your answer \[y-f(3)=f'(3)(x-3)\]

OpenStudy (anonymous):

yw